OpenStudy (anonymous):

find the directional derivative of f at the point P in the indicated direction: f(xy) = xcos^2y p(2, pi/4) < 5, 1>

OpenStudy (anonymous):

i just want to make sure i have my Fx and Fy correct.. i always have difficulties with cos/sin Fx=cos^2y Fy= -2xsin^2y

OpenStudy (kinggeorge):

\(f_x\) looks correct, but \[f_y=-2x \cos(y)\sin^2(y)\]

OpenStudy (anonymous):

how come? :(

OpenStudy (kinggeorge):

\[{\partial \over \partial y}\;\; x \cos^2(y)=2x \cos(y) \cdot \left( {\partial \over \partial y} \cos(y)\right)=-2x \cos(y) \sin(y)\]I apologize. I shouldn't have squared sin.

OpenStudy (anonymous):

ohhhhh okay! cuz i have to derive the cos, that's right.. okay, cool, thank you!